# Picosecond Wireless Synchronization with Entangled Photons via Grid-Based Quantum Coverage in Indoor Optical Systems

### v1

@@ Abstract
In this paper, we propose a novel entanglement-assisted synchronization framework for indoor optical wireless systems based on grid-based beam steering. A central transmitter equipped with a spontaneous parametric down-conversion (SPDC) source emits time-energy entangled photon pairs, directing user photons toward spatially defined grids based on estimated user positions, while retaining reference photons for timestamping. The room is partitioned into multiple beam-aligned regions, and photon reception probability is analytically modeled using local coordinate transformations and Gaussian beam optics. To address the inherent sparsity and randomness of photon detection, we develop a robust two-stage synchronization algorithm that performs sparse bit-pattern matching and timestamp averaging to estimate timing offsets. Monte Carlo simulations are then performed to evaluate the impact of synchronization duration, grid resolution, and photon-pair generation rate on synchronization accuracy. Results show that finer grid configurations and optimal photon-pair rates significantly reduce synchronization error, achieving sub-$10$ ps timing accuracy within $1\,\text{ms}$ of synchronization time. The proposed approach enables scalable, high-precision quantum synchronization in wireless indoor environments, laying the groundwork for future joint quantum communication, sensing, and positioning systems.
@@ I Introduction
The convergence of quantum information science and wireless communication is driving the development of a new class of systems that promise secure communication, high-precision sensing, and distributed quantum computing [1, 2]. Among these, optical wireless communication (OWC) systems, which use light for high-speed data transfer, have emerged as a promising platform for enabling quantum-enhanced applications, including quantum key distribution (QKD), quantum sensing, and local-area quantum networks [2, 3, 4, 5, 6]. A fundamental requirement across nearly all such applications is the ability to distribute and detect entangled or single photons with extremely high temporal precision, often at the picosecond scale [7].
@@
This stringent requirement for temporal precision becomes especially critical in protocols such as QKD and entanglement swapping, where synchronization between a central node and multiple users must reach the picosecond level. In these scenarios, the identification of correlated photon detection events depends on extremely narrow coincidence windows. Even minor clock offsets or uncertainties in propagation delay can significantly degrade system performance, increasing error rates and undermining quantum coherence [8, 9]. As a result, time synchronization is not merely an engineering challenge but a fundamental physical necessity in quantum communication systems.
@@
Given the critical role of synchronization in quantum systems, the field of quantum clock synchronization (QCS) has evolved along two primary paths. The first focuses on long-haul fiber-optic quantum links, where weak quantum signals must coexist with classical data traffic [10, 11]. These systems employ techniques such as wavelength division multiplexing and polarization management to mitigate crosstalk and dispersion effects. Recent advances have further extended this, where quantum-enhanced two-way time transfer leverages the strong temporal correlations in energy-time entangled photons to achieve precise synchronization stability [12, 13]. The second path targets long-distance free-space quantum links, including ground-to-satellite communication, where atmospheric turbulence, solar background radiation, and Doppler shifts pose significant challenges. Solutions in this domain typically involve adaptive optics, narrowband filtering, time transfer, and probabilistic error correction [14, 15, 16]. While both approaches are technically sophisticated, their assumptions and mitigation strategies are highly specialized and not directly applicable to indoor environments.
While existing QCS approaches have focused on long-haul fiber or free-space links, a parallel line of development in the classical domain has explored scalable indoor OWC architectures. Among these, the Grid-of-Beams (GoB) approach has emerged as a practical solution for supporting mobile users [17, 18, 19, 20, 21]. In a GoB system, a dense array of static, narrow laser beams provides full spatial coverage, and user mobility is managed through digital handovers between adjacent grid cells, thereby eliminating the need for continuous beam tracking [17, 19]. However, these systems are inherently quantum-agnostic. Each handover alters the optical path length, disrupting prior timing calibration. Moreover, current GoB architectures lack mechanisms to rapidly re-establish the picosecond-level synchronization required by quantum protocols. More broadly, existing QCS schemes assume stable links and long integration times, making them unsuitable for dynamic indoor environments characterized by frequent handovers and ambient lighting noise. This highlights a critical gap at the intersection of quantum synchronization and mobile indoor OWC that must be addressed to enable practical quantum networking in real-world settings.
To bridge this gap, we propose the first integrated framework for entanglement-assisted picosecond synchronization tailored to mobile indoor optical wireless environments. Our approach leverages the temporal correlations of entangled photon pairs in conjunction with a scalable optical delivery system to enable precise synchronization. This method allows for accurate estimation of timing parameters without relying on a shared clock or fixed spatial configuration, making it well-suited for dynamic, mobile scenarios.
To enable practical deployment in dynamic indoor environments, we build upon existing and experimentally validated GoB architectures, originally developed for classical OWC systems to provide full indoor coverage and support user mobility [17, 22]. The GoB discretizes space into fixed beams, each targeting a predefined coverage cell, thereby simplifying link acquisition and handover. We extend this architecture by embedding a quantum-native synchronization layer, enabling support for entanglement-based protocols. We complement this architecture with a two-stage synchronization algorithm designed to operate over sparse and noisy single-photon detection streams. The first stage performs coarse alignment using bit-pattern matching, while the second refines synchronization through timestamp averaging. To support system design and performance evaluation, we also develop a comprehensive analytical model that captures photon-pair generation statistics, Gaussian beam propagation, user position uncertainty, and timing jitter from both the source and detectors.
To the best of our knowledge, this is the first work to present and analyze in detail an entanglement-assisted synchronization framework tailored to a grid-based indoor OWC system. Our main contributions are as follows:
  • We propose a complete system architecture for high-precision timing in indoor environments, integrating a grid-based optical delivery system with an entanglement-based synchronization protocol. This design directly addresses the challenges of user mobility and line-of-sight link maintenance.
  • We develop a detailed, end-to-end analytical model that captures the key physical phenomena. This includes the Poisson statistics of the quantum unit, Gaussian beam optics for photon propagation, the impact of user positioning errors on reception probability, and a rigorous treatment of both source and detector timing jitter.
  • We design a practical algorithm specifically to handle the sparse and noisy data characteristic of single-photon channels. The algorithm performs sparse bit-pattern matching for coarse slot-level alignment, followed by timestamp averaging over high-confidence matched events to achieve fine-grained, picosecond-level precision.
  • Through Monte Carlo simulations, we evaluate the performance of our framework under realistic conditions. We demonstrate that sub-$10$ picosecond synchronization accuracy is achievable within a $1$ ms timeframe. Our results quantify the significant impact of spatial grid resolution and reveal the existence of an optimal photon-pair generation rate that balances signal availability against multi-photon ambiguity.
@@
Thus, this work provides a foundational step toward building practical, scalable, and high-precision quantum networks in local indoor settings, bridging quantum timing principles with proven classical networking architectures.
@@
The remainder of this paper is structured as follows. Section II presents the system model considered in this study. Section III describes the spatial modelling of photon reception in the communication setup. Section IV introduces the proposed synchronization models. Simulation results are discussed in Section V. Finally, Section VI concludes the paper and outlines directions for future research.
TABLE I: List of Key Notations
@@ II System Model
We consider an indoor quantum synchronization system based on a grid-structured optical wireless delivery architecture that utilizes time-energy entangled photon pairs. A central quantum unit, mounted on the ceiling, contains a spontaneous parametric down-conversion (SPDC) source that continuously generates entangled photons. In each pair, one photon remains at the central unit as a precise timing reference, while the other is directed toward the user through one of $N_{g}$ spatially distinct optical wireless transmission paths. Each transmission path corresponds to a fixed illumination region (grid) on the indoor surface, effectively dividing the room into $N_{g}$ non-overlapping zones. In the following subsections, we provide a detailed description of the system model and the parameters considered in this paper. Meanwhile, a summary of the notations used is presented in Table I.
Figure 1: System architecture of the proposed entanglement-based indoor quantum synchronization scheme. A central ceiling-mounted transmitter emits time-energy entangled photon pairs, directing the user photon toward a spatially selected beam corresponding to the estimated user location. The beam is aligned with a specific grid region in the room, while the reference photon is time-stamped at the transmitter. The receiver structure includes a circular aperture and SPAD array aligned to detect the photon and record its arrival time for synchronization. Local $(x^{\prime},y^{\prime},z^{\prime})$ and global $(x,y,z)$ coordinate systems are depicted for spatial modeling.
@@ II-A Geometric Configuration and Coordinate Framework
Let $\mathbf{p}_{u}=[x_{u},y_{u},z_{u}]$ denote the true position of the user, and let $\hat{\mathbf{p}}_{u}=[\hat{x}_{u},\hat{y}_{u},\hat{z}_{u}]$ represent the estimated position available to the system. The estimation is subject to random errors modeled as independent Gaussian noise in each coordinate. Specifically, we have
@@ II-B Transmitter Architecture and Photon Pair Emission
The central transmitter unit employs a SPDC source to generate time-energy entangled photon pairs. Each generated pair consists of two photons: one designated as the reference photon and the other as the user photon. The reference photon is routed to a time-stamping detection module co-located at the transmitter, while the user photon is directed toward a spatially-resolved optical output based on the user’s estimated position, as shown in Fig. 1.
@@ II-C Receiver Architecture and Photon Capture
where $t_{a}$ is the actual photon arrival time, and $t_{d}$ is the recorded detection time. The value of the standard deviation parameter, $\sigma_{d}$, depends on the SPAD’s material and circuit design, and typically ranges from $20$ ps to $100$ ps in modern systems [30, 31].
@@ IV-C Synchronization Recovery at the User
V Simulations

### v2

@@ Abstract
In this paper, we present a robust entanglement-assisted synchronization framework for indoor optical wireless systems that explicitly captures the coupling between spatial beam geometry and temporal synchronization accuracy. Unlike conventional approaches that treat beam steering and timing estimation independently, a unified spatio temporal model is developed that links user position uncertainty to the Cramér Rao lower bound of the synchronization error. The framework incorporates key physical impairments, including multipath dispersion, non Gaussian detector jitter, and spatially correlated localization errors. Through analytical modeling and extensive simulations, we show that the proposed system exhibits graceful performance degradation under heavy tailed positioning uncertainty and remains stable in the presence of multipath induced bias. Using realistic single photon detector parameters, the results indicate that synchronization accuracy below $10$ picoseconds can be maintained across a wide range of operating conditions. This level of precision provides a scalable foundation for quantum enabled indoor wireless networks.
@@ I Introduction
The convergence of quantum information science and wireless communication is enabling a new class of systems that support secure communication, high precision sensing, and distributed quantum processing [1, 2]. Among the available physical platforms, optical wireless communication has emerged as a particularly attractive candidate for realizing quantum enhanced networking functionalities, including quantum key distribution, quantum sensing, and local area quantum networks [2, 3, 4, 5, 6]. A fundamental requirement across all such applications is the ability to distribute and detect entangled or single photons with extremely high temporal precision, typically at the picosecond level [7].
@@
This requirement becomes especially critical in protocols such as quantum key distribution and entanglement swapping, where the identification of correlated detection events relies on narrow coincidence windows. Even small clock offsets or uncertainties in propagation delay can significantly degrade detection fidelity, increase error rates, and compromise quantum coherence [8, 9]. Consequently, time synchronization is not merely an implementation detail but a core physical requirement for reliable quantum communication.
TABLE I: System-Level Comparison of Representative Synchronization Schemes
Feature Classical Two-Way Single-Beam Proposed
Optical Quantum Quantum Grid-Quantum
Spatial Coverage No No No Yes
Mobility Support Limited No Limited Yes
Sync Time Fast Long Moderate Short
Robust to Sparsity No Moderate Low Yes
Dominant Noise Thermal Fiber Jitter Shot Noise Shot Noise/Multipath
Sensitivity Poor (High power required) Medium Good Good
Link Model Continuous Static P2P Continuous Grid-Based
Handover Support Complex (Re-locking) No No Native (Fast Re-sync)
Note: Sensitivity and synchronization times are indicative, corresponding to different operating regimes. Classical schemes assume continuous high-SNR detection; quantum and SPAD schemes operate in photon-counting regimes with distinct noise statistics.
@@
Motivated by this need, research on quantum clock synchronization has progressed along two principal directions. The first focuses on long haul fiber based quantum links, where weak quantum signals must coexist with classical traffic [10, 11]. These systems rely on techniques such as wavelength division multiplexing, polarization control, and dispersion compensation, and have recently demonstrated high precision synchronization using energy time entangled photons [12, 13]. The second direction targets free space quantum links, including ground to satellite communication, where synchronization must contend with atmospheric turbulence, background radiation, and Doppler effects [14, 15, 16]. While both approaches have achieved impressive performance, they are designed for relatively static links and rely on assumptions that do not hold in indoor environments.
Indoor optical wireless systems operate under conditions that differ fundamentally from long-haul or point-to-point optical links. They must support user mobility, frequent handovers, dense spatial reuse, and operation in multipath-rich environments with ambient optical interference. OWC is well suited to these conditions due to its large available bandwidth, strong spatial confinement, and compatibility with single-photon detection technologies. However, these same characteristics impose stringent requirements on synchronization, as timing accuracy must be maintained despite spatial uncertainty and dynamic link reconfiguration. Consequently, indoor quantum synchronization schemes must be inherently robust to user motion and beam misalignment, rather than relying on static or continuously aligned links.
Despite recent progress in entanglement assisted synchronization and optical time transfer [17], most existing formulations implicitly decouple spatial beam configuration from timing performance. Synchronization accuracy is typically analyzed under fixed link assumptions, without accounting for the effects of beam footprint, user position uncertainty, or spatial coverage geometry on photon detection statistics. As a result, the interaction between spatial design and achievable timing precision remains largely unexplored, even though it is central to practical indoor optical wireless deployments.
At the same time, significant progress has been made in classical indoor optical wireless architectures. In particular, the grid of beams (GoB) paradigm has emerged as an effective solution for providing full spatial coverage and supporting user mobility [18, 19, 20, 21, 22]. In a GoB system, a dense array of static, narrow beams partitions the coverage area into fixed cells, and user mobility is handled through digital handovers rather than mechanical beam steering. Although this architecture offers strong scalability and robustness, it is inherently quantum-agnostic. Each handover modifies the optical path length and invalidates prior timing calibration, while existing GoB implementations lack mechanisms to rapidly re-establish synchronization at the precision required for quantum applications under user mobility.
In this work, we address this gap by developing a unified spatio temporal synchronization framework in which spatial grid geometry and timing performance are analytically coupled. Unlike prior approaches that treat beam steering and synchronization independently, we show that user position uncertainty directly determines the number of detected photon pairs and therefore the fundamental synchronization accuracy. This relationship is captured through an explicit Cramér Rao lower bound (CRLB) formulation. Thus, it enables quantitative evaluation of how spatial design choices translate into timing precision. To clarify the limitations of existing approaches and to position the proposed method within the broader landscape, Table I presents a system level comparison of representative synchronization schemes. The comparison highlights that existing classical and quantum approaches either lack spatial scalability or fail to support mobility and sparse photon detection, motivating the need for a grid based quantum synchronization framework.
Building on this insight, we particularly propose an entanglement assisted synchronization framework specifically tailored to grid based indoor optical wireless systems. The proposed approach combines the spatial scalability of GoB architectures with the temporal precision offered by entangled photon pairs. A two stage synchronization algorithm is then introduced, consisting of coarse alignment using sparse bit pattern matching followed by fine synchronization through timestamp averaging over correlated detection events. In parallel, a comprehensive analytical model is developed to capture photon pair statistics, Gaussian beam propagation, user position uncertainty, and timing jitter introduced by both the source and detectors. To the best of our knowledge, this is the first work to present a complete entanglement assisted synchronization framework explicitly designed for grid based indoor optical wireless networks.
The main contributions of this paper are summarized as follows:
  • We introduce a unified system architecture that integrates grid based optical delivery with entanglement assisted synchronization. This design establishes the fundamental coupling between spatial beam footprint and achievable timing variance.
  • Next, we develop an end to end analytical model that extends beyond ideal assumptions to incorporate multipath induced dispersion, non Gaussian positioning errors, and heavy tailed detector jitter.
  • Also, we design a practical two stage synchronization algorithm tailored for sparse single photon detection and robust to the timing biases and spatial misalignments inherent to indoor mobility.
  • Finally, we conduct a comprehensive performance evaluation demonstrating sub 10 ps accuracy and verifying system stability under realistic non Gaussian error distributions.
@@
This work therefore provides a foundational step toward scalable and high precision quantum networking in indoor environments, bridging quantum synchronization theory with practical optical wireless system design.
@@
The remainder of this paper is organized as follows. Section II presents the system model. Section III describes the spatial modeling of photon reception. Section IV introduces the synchronization framework and the associated theoretical analysis. Section V derives the CRLB and establishes fundamental limits on synchronization accuracy. Section VI introduces the channel impairments considered in the proposed framework. Simulation results and performance evaluation are presented in Section VII. Finally, Section VIII concludes the paper and outlines directions for future research.
TABLE II: List of Key Notations
@@ II System Model
We consider an indoor quantum synchronization system based on a grid-structured optical wireless delivery architecture that utilizes polarization entangled photon pairs. A central quantum unit, mounted on the ceiling, contains a spontaneous parametric down-conversion (SPDC) source that continuously generates entangled photons. In each pair, one photon remains at the central unit as a precise timing reference, while the other is directed toward the user through one of $N_{g}$ spatially distinct optical wireless transmission paths. Each transmission path corresponds to a fixed illumination region (grid) on the indoor surface, effectively dividing the room into $N_{g}$ non-overlapping zones. In the following subsections, we provide a detailed description of the system model and the parameters considered in this paper. Meanwhile, a summary of the notations used is presented in Table II.
Figure 1: System architecture of the proposed entanglement-based indoor quantum synchronization scheme. A central ceiling-mounted transmitter emits polarization encoded entangled photon pairs, directing the user photon toward a spatially selected beam corresponding to the estimated user location. The beam is aligned with a specific grid region in the room, while the reference photon is time-stamped at the transmitter. The receiver structure includes a circular aperture and SPAD array aligned to detect the photon and record its arrival time for synchronization. Local $(x^{\prime},y^{\prime},z^{\prime})$ and global $(x,y,z)$ coordinate systems are depicted for spatial modeling.
@@ II-A Geometric Configuration and Coordinate Framework
Let $\mathbf{p}_{u}=[x_{u},y_{u},z_{u}]$ denote the true position of the user, and let $\hat{\mathbf{p}}_{u}=[\hat{x}_{u},\hat{y}_{u},\hat{z}_{u}]$ represent the estimated position available to the system. Here, we assume that the estimation is subject to random errors modeled as independent Gaussian noise in each coordinate. However, in the following subsections we investigate how alternative error models affect the performance of the synchronization system. Accordingly, we have
@@ II-B Transmitter Architecture and Photon Pair Emission
The central transmitter unit employs a SPDC source to generate polarization encoded entangled photon pairs. Each generated pair consists of two photons: one designated as the reference photon and the other as the user photon. The reference photon is routed to a time-stamping detection module co-located at the transmitter, while the user photon is directed toward a spatially-resolved optical output based on the user’s estimated position, as shown in Fig. 1.
@@ II-C Receiver Architecture and Photon Capture
where $t_{a}$ is the actual photon arrival time, and $t_{d}$ is the recorded detection time. The value of the standard deviation parameter, $\sigma_{d}$, depends on the SPAD’s material and circuit design, and typically ranges from $20$ ps to $100$ ps in modern systems [30, 31]. The effective number of usable photon pairs available for synchronization is therefore jointly determined by the reception probability and the timestamp quality, which motivates the coupled spatial and temporal analysis developed in the following sections.
@@ III Spatial Modeling of Photon Reception via Aperture
It is also worth noting that, although independent Gaussian perturbations for spatial uncertainty are used as the baseline model for analytical tractability, the framework is not limited to Gaussian localization errors. In later sections, we also examine heavy tailed, biased, and spatially correlated position errors with matched root mean square deviation to assess robustness and to isolate the influence of spatial uncertainty on photon reception statistics.
@@ IV-C Synchronization Recovery at the User
V Fundamental Limits: Spatio-Temporal Cramér-Rao Bound
To establish the fundamental performance limits of our synchronization framework and demonstrate the novel spatio-temporal coupling introduced by the grid architecture, we derive the CRLB for the synchronization offset estimate $\hat{\tau}_{\text{sync}}$. This analysis reveals how spatial grid parameters directly influence the achievable timing precision, a distinctive feature of our quantum wireless system that has no classical counterpart.
V-1 Likelihood Function for Timestamp Differences
After successful bit-pattern matching (Algorithm 1), we obtain a set $\mathcal{J}_{\text{match}}$ of reliably matched slots with cardinality $M=|\mathcal{J}_{\text{match}}|$. For each matched slot $j$, we observe the timestamp difference as
$\Delta_{j}=t_{u}[i_{j}]-t_{r}[i_{j}-\ell^{*}],$ (31)
where $t_{r}[k]$ and $t_{u}[k]$ are given by (18) and (24), respectively. The reference timestamp is for a successfully detected photon pair. Accordingly, one can rewrite (31) as
$\Delta_{j}=\tau_{\text{sync}}+(n_{u}[i_{j}]-n_{r}[i_{j}-\ell^{*}]).$ (32)
Given the independence of $n_{u}$ and $n_{r}$, the difference in (32) follows
$\Delta_{j}\sim\mathcal{N}(\tau_{\text{sync}},2\sigma_{d}^{2}).$ (33)
The likelihood function for $M$ independent observations can thus be written as
$f(\{\Delta_{j}\}_{j=1}^{M}|\tau_{\text{sync}})=\prod_{j=1}^{M}\frac{1}{\sqrt{4\pi\sigma_{d}^{2}}}\exp\left(-\frac{(\Delta_{j}-\tau_{\text{sync}})^{2}}{4\sigma_{d}^{2}}\right).$ (34)
V-2 Fisher Information and CRLB
Let’s write the log-likelihood function with respect to the synchronization delay parameter $\tau_{\text{sync}}$ as
$\mathcal{L}(\tau_{\text{sync}})=-\frac{M}{2}\ln(4\pi\sigma_{d}^{2})-\sum_{j=1}^{M}\frac{(\Delta_{j}-\tau_{\text{sync}})^{2}}{4\sigma_{d}^{2}}.$ (35)
The Fisher Information for $\tau_{\text{sync}}$ is then obtained as
$I(\tau_{\text{sync}})=-\mathbb{E}\left[\frac{\partial^{2}\mathcal{L}}{\partial\tau_{\text{sync}}^{2}}\right]=\sum_{j=1}^{M}\frac{1}{2\sigma_{d}^{2}}=\frac{M}{2\sigma_{d}^{2}}.$ (36)
Thus, the CRLB for any unbiased estimator $\hat{\tau}_{\text{sync}}$ is
$\text{Var}(\hat{\tau}_{\text{sync}})\geq\frac{1}{I(\tau_{\text{sync}})}=\frac{2\sigma_{d}^{2}}{M}.$} (37)
This bound holds for any unbiased estimator conditioned on $M$ successfully matched photon pairs and shows that the achievable timing precision scales inversely with the number of reliable detections.
V-3 Spatial Dependence of Matched Pair Count
The number of matched pairs $M$ is a random variable whose expectation depends critically on spatial parameters. More precisely, for a synchronization sequence comprising $N_{s}=T_{\text{seq}}/T_{\text{qs}}$ slots, a valid single‑pair emission occurs with probability $\mu_{t}e^{-\mu_{t}}$ under Poisson statistics. Conditioning on a single‑pair event, a successful match further requires photon reception, detection, and the absence of background‑induced ambiguity. Hence, the expected number is
$\mathbb{E}[M]=N_{s}\cdot\mu_{t}e^{-\mu_{t}}\cdot\overline{P}_{\text{match}}(N_{g},\sigma_{p}),$ (38)
where $\overline{P}_{\text{match}}(N_{g},\sigma_{p})$ is the average probability that a generated photon pair results in a successful match. This matching probability incorporates spatial alignment effects through
$\overline{P}_{\text{match}}(N_{g},\sigma_{p})=\eta_{d}(1-p_{\text{back}})\cdot\mathbb{E}[P_{\text{recept}}^{(i,j)}],$ (39)
where $\eta_{d}$ is the detector efficiency, $p_{\text{back}}$ accounts for background-induced failures, i.e., the probability that background photons cause a slot to be invalid for synchronization. Also, $\mathbb{E}[P_{\text{recept}}^{(i,j)}]$ is the expected reception probability over user position uncertainty (using the spatial reception model developed in Section III). From (16), the reception probability for a beam aligned with grid $(i,j)$ is
$P_{\text{recept}}^{(i,j)}\approx\frac{2r_{a}^{2}}{w^{2}(z_{u}^{\prime})}\exp\left(-\frac{2[(x_{u}^{\prime})^{2}+(y_{u}^{\prime})^{2}]}{w^{2}(z_{u}^{\prime})}\right),$ (40)
where $(x_{u}^{\prime},y_{u}^{\prime})$ represents the user’s offset from the beam center in local coordinates. With Gaussian position uncertainty $\sigma_{p}^{2}$ in each coordinate, the expected reception probability becomes
$\mathbb{E}[P_{\text{recept}}^{(i,j)}]=\frac{2r_{a}^{2}}{w^{2}(z_{u}^{\prime})+4\sigma_{p}^{2}}.$ (41)
Finally, combining (37), (41), and (38), we obtain the CRLB that yields the following information-theoretic bound
$\mathrm{Var}(\hat{\tau}_{\mathrm{sync}})\geq\frac{\sigma_{d}^{2}}{N_{s}\mu_{t}e^{-\mu_{t}}\eta_{d}\left(1-p_{\mathrm{back}}\right)}\cdot\frac{w^{2}(z_{u}^{\prime})+4\sigma_{p}^{2}}{r_{a}^{2}}.$ (42)
This bound reveals a direct coupling between spatial beam design and temporal synchronization accuracy. In particular, reducing the beam width $w(z_{u}^{\prime})$ through finer grid resolution improves timing precision only until $w^{2}(z_{u}^{\prime})$ becomes comparable to the position‑uncertainty term $4\sigma_{p}^{2}$. Beyond this point, further grid refinement yields diminishing returns, as synchronization performance becomes dominated by localization uncertainty rather than beam focusing.
This behavior highlights a fundamental distinction between classical and quantum grid‑based systems. While classical GoB architectures primarily trade spatial resolution for received optical power, in quantum synchronization systems the spatial architecture explicitly enters the information bound governing timing accuracy. To the best of our knowledge, this spatio‑temporal coupling has not been previously characterized in the context of indoor entanglement‑assisted synchronization.
VI Channel Impairments and Timing Uncertainty Model
In the previous sections, idealized channel conditions were assumed in order to isolate the fundamental operation of the synchronization framework. In this section, we introduce the channel impairments considered in the proposed framework and formalize their impact on the timing observations used for synchronization. The purpose of this section is to define the physical mechanisms that degrade timing accuracy in indoor optical wireless environments, independently of the specific synchronization algorithm.
VI-A Multipath Propagation and Non-Line-of-Sight Components
In indoor optical wireless systems, photons can reach the receiver through both a direct line-of-sight (LOS) path and multiple reflected paths arising from walls, ceilings, and surrounding objects. While the LOS component carries the correct timing reference, reflected components arrive with excess delay and introduce temporal dispersion. To model this effect, the optical channel impulse response is represented using a double-exponential form commonly adopted for indoor optical wireless channels [35, 36, 37]:
$h_{\text{total}}(t)=\eta_{\text{LOS}}\delta(t)+\sum_{k=1}^{N_{\text{ref}}}\eta_{k}\exp\!\left(-\frac{t-\Delta t_{k}}{\tau_{\text{rms}}}\right),$ (43)
where $\eta_{\text{LOS}}$ denotes the weight of the LOS component, $\eta_{k}$ represents the contribution of the $k$-th reflected path, $\Delta t_{k}$ is its associated excess delay, and $\tau_{\text{rms}}$ is the root-mean-square (RMS) delay spread of the channel.
Here, we assume that the LOS term dominates in typical indoor environments; however, reflected components introduce a non-negligible tail in the impulse response [36]. We also assume the RMS delay spread is set to $\tau_{\text{rms}}=300$ ps, consistent with reported indoor optical measurements [38]. This multipath induced dispersion introduces a systematic bias in the detected photon arrival times and represents a fundamental limitation on synchronization accuracy, which we examine in detail in the performance evaluation section.
VI-B Non-Gaussian Timing Jitter and Diffusion Effects
In addition to channel induced dispersion, the temporal resolution of the receiver is constrained by the intrinsic timing jitter of the SPAD. A Gaussian model was used in the previous section for analytical tractability. However, although Gaussian jitter models are commonly adopted for convenience, experimental studies have shown that SPAD timing responses exhibit pronounced long tailed behavior caused by carrier diffusion within the absorption and multiplication regions of the detector [29].
To capture this effect and to assess the robustness of the proposed synchronization framework under realistic conditions, we also model the detector timing jitter using a Laplacian distribution and provide the performance evaluation. This choice is physically justified by the exponential nature of diffusion-induced delays and has been widely adopted in photon-timing experiments [39, 40]. The corresponding probability density function is given by
$f(x\mid\mu,b)=\frac{1}{2b}\exp\!\left(-\frac{|x-\mu|}{b}\right),$ (44)
where $\mu$ denotes the nominal arrival time and the scale parameter $b$ is related to the standard deviation $\sigma$ via $b=\sigma/\sqrt{2}$.
The Laplacian model exhibits heavier tails than a Gaussian distribution and therefore provides a conservative representation of timing uncertainty. To have a fair comparison, in this work, the Laplacian parameters are selected such that the variance matches that of the Gaussian jitter model used in the baseline case in the previous section.
VI-C Unified Timing Observation Model
Taking into account photon generation uncertainty, multipath dispersion, and detector timing jitter, the timestamp recorded at the receiver (i.e., (24)) can be expressed as
$t_{u}=t_{g}+\tau_{\mathrm{sync}}+\Delta t_{\mathrm{multipath}}+n_{\mathrm{jitter}},$ (45)
where $t_{g}$ is the photon generation time, $\tau_{\mathrm{sync}}$ denotes the clock offset between the transmitter and receiver, $\Delta t_{\mathrm{multipath}}$ represents the excess delay caused by reflected paths, and $n_{\mathrm{jitter}}$ follows the Laplacian distribution defined above. This formulation explicitly separates deterministic synchronization offsets from stochastic channel and detector effects. Importantly, it highlights that multipath propagation introduces a bias term that does not vanish with increasing photon count, while detector jitter determines the convergence behavior of the synchronization estimator. These effects are examined quantitatively in the performance evaluation section, where their impact on synchronization accuracy and convergence behavior is assessed.
VII Simulations